# Deep Emotion Recognition in Textual Conversations: A Survey

### v3

@@ 3.1.3 Mixed Emotions
A person can experience simultaneuosly several emotions. For example, one can experience joy and surprise at the same time and sometimes even opposite emotions simultaneously [4]. While the VAD model allows for an unlimited description of emotions so that a mixture of emotions can be expressed in a single VAD description, the single label categorical approach does not account for mixed emotions. This motivates the need to classify emotions with multiple labels when using the categorical approach. This need is very obvious when using a low number of labels, but there are also multi-label datasets with a high number of emotion labels, such as the GoEmotions dataset [18].
@@ 3.1.4 ERC Taxonomy Considerations
The existent taxonomies are usually transversally applied along all Emotion Recognition tasks. However, for ERC, it can be observed that there are coincident predominant classes in several benchmark datasets, such as Neutral and Happiness/Joy, as it can be seen on the class distribution of the datasets on subsection 3.2. Within different ERC domains such as chit-chat vs task-oriented, such as mental health [21] and customer support [39], there are specific classes of emotions that will be more represented than others in each domain. With such a rich diversity of possible classes for this task, the dimensional approach can encompass all classes but might make it harder to distinguish between differentiated emotional displays.
@@ 3.2 ERC Benchmark Datasets
In this subsection, different categorical, single label datasets are presented, that were used by the works benchmarked in this survey. These works are organized on Tables LABEL:ercmul and 8. Other relevant datasets are also presented. Overall, their class imbalance is significant suggesting the use of balancing techniques that will be elaborated on in subsection 5.2. All datasets are in English, which is a limitation concerning multilingual ERC, as pointed out in section 2.
@@ 3.2.1 IEMOCAP [8]
The IEMOCAP dataset consists in 151 videos of two speakers’ dialogues, comprising 7.433 utterances with an average length of 11,6 words. Each utterrance is labelled with one of 8 emotions: happy, sad, neutral, angry, excited, frustrated fear and disgust. Most of the works, however, benchmark their performance on the first 6 classes, whose label distribution is displayed on Table 1. The maximum class imbalance ratio is 1:3 (happy:frustrated).
Table 1: IEMOCAP dataset label distribution.
Hap Sad Neutral Ang Exc Frust
648 1084 1708 1103 1041 1849
Sample: "Is that, is that - is that just foam? I can’t even tell. Although, if you can’t tell, it’s probably isn’t them. It’ll probably be unmistakable, don’t you think?" Label: Excited
@@ 3.2.2 MELD [97]
The Multimodal EmotionLines Dataset (MELD) contains about 13.000 utterances from 1.433 dialogues from the TV-series Friends. Each utterance encompasses audio, visual, and textual modalities, with average length of 8,0 words. Emotion categories are Ekman’s basic emotions and neutral. The label distribution is given in Table 2. The maximum class imbalance ratio is 1:18 (fear:neutral).
Table 2: Friends dataset label distribution.
Neutral Ang Dis Fea Joy Sad Sur
6436 1607 361 358 2308 1002 1636
Sample: "Are we okay now?" Label: Fear
@@ 3.2.3 EmoryNLP [139]
EmoryNLP is a text dataset extracted from transcripts of the TV show Friends. It contains 12.606 utterances, labeled with one of the six primary emotions in Willcox’s feeling wheel [127]: sad, mad, scared, powerful, peaceful, joyful, and a neutral label, whose distribution in the dataset is displayed on Table 3. The maximum class imbalance ration is 1:4 (sad:neutral).
Table 3: EmoryNLP dataset label distribution.
Neutral Joy Peace Pow Scared Mad Sad
3776 2755 1190 1063 1646 1332 844
Sample: "What a coincidence, I listen in my sleep." Label: Peaceful
@@ 3.2.4 DailyDialog [70]
DailyDialog is a large text dataset built from websites used to practice English dialogue in daily life, containing 12.218 multi-turn dialogues, comprising 102.879 utterances with an average length of 14,6 words. Each utterance is labelled with one of the six Ekman’s basic emotions or other. The label distribution is displayed on table 4. The maximum class imbalance ratio is 1:1156 (fear:other).
Table 4: DailyDialog dataset label distribution.
Ang Dis Fear Hap Sad Sur Other
1022 353 74 12885 1150 1823 85572
Sample: "I was scared stiff of giving my first performance" Label: Disgust
3.2.5 Friends [42]
Friends consists in a Friends TV Scripts corpus in which a scene is treated as a dialogue, containing 1000 text dialogues and 17.799 utterances with an average length of 10,7 words. Each utterance is labelled with one of the six Ekman’s basic emotions and neutral and non-neutral, assigned when the utterance has more than two emotions. The label distribution is given in Table 5. The maximum class imbalance ratio is 1:11 (fear:non-neutral).
Table 5: Friends dataset label distribution.
Neutral Joy Sad Fea Ang Sur Dis Non-neu
7565 2215 619 283 900 2104 389 3724
Sample: "Oh my God, oh my God! Poor Monica!" Label: Surprise
3.2.6 EmoContext [10]
EmoContext consists in textual dialogues from user interaction with a conversational agent. It contains 38.424 dialogues with three turns each, with an average lenght of 7,6 words per turn, and the dialogues contain emojis. Each dialogue is labeled as angry, happy, sad or others, whose label distribution is displayed on Table 6. The maximum class imbalance ratio is 1:5 (happy:others).
Table 6: EmoContext dataset label distribution.
Happy Sad Angry Others
4669 5838 5954 21963
Sample: "How dare you to slap my child If you spoil my car, I will do that to you too Just try to do that once" Label: Angry
@@ 3.2.7 Other Datasets
The RECOLA dataset [103] is a multimodal corpus of spontaneous human interactions collected from a collaborative task. It includes not only audio and video, but also electrocardiogram and electrodermal activity data. Emotional behaviours are annotated in terms of arousal and valence, in continuous time, and social behaviours in terms of agreement, dominance, engagement, performance and rapport, in discrete time.
@@ 4 Deep Learning for Emotion Recognition in Conversations
In this section, key ERC works are described, focusing on the challenges and opportunities of the conversational scenario, presented in Section 2 and on the diverse Deep Learning architectures that can be employed for the task. Descriptions of key Deep Learning models are provided along with descriptions of works that resorted to those models. The works on Emotion Recognition in Conversations using Deep Learning described in this survey were chosen to be representative of the different techniques that are employed, covering gated Neural Networks with and without bidirectionality, Memory Networks, Graph Neural Networks, Attention Mechanisms and Transformers. It is remarkable to observe how in four years research evolved form a simple LSTM, taking into account the context of the conversation, to complex gated and graph neural network architectures and how the invention of transformers was reflected in not only word embeddings, such as BERT, but also in new classifier architectures. Comparisons between these works are performed based on the reported weighted-F1-score without the majority class on the IEMOCAP dataset, the choice of most authors including the prominent works of [95], [37], [78], [31], [149], [107], [30], [122], [65], [108], [108] and [62], since the dataset is not too imbalanced.
@@ 4.1 Convolutional Neural Networks for Feature Extraction
Convolutional Neural Networks (CNNs) are a type of feedforward neural networks inspired by the biology of the visual cortex that mimic the local filtering over the input space performed by the receptive fields. In the case of text, a linear layer is applied to a sequence of words, each represented by its embeddings, separately and in the same way for each sequence. The convolutional layer can have many filters, each extracting a different feature, and is usually followed by a pooling operation. In the case of Emotion Recognition in Conversations, convolutional layers act as feature extractors, useful since strong predictors of the label can appear in different places in the input. With the advent of Pre-Trained Language Models for utterance representation, elaborated on subsubsection 4.9.4, convolutional neural networks have been less used.
4.2 The Multi Layer Perceptron
Emotion Recognition is a classification task. The most simple neural network architecture that can be employed for such task is the Multi Layer Perceptron.
The Multi Layer Perceptron (MLP) is a non-linear function approximator that can be used for classification and regression. It is a class of feed-forward artificial neural networks that is composed of at least three layers: an input layer, a hidden layer and an output layer. Its complexity and representation power increases with the number of layers. A MLP is mathematically described by the following equation:
@@ 4.2 The Multi Layer Perceptron
In which L is the number of layers, $W_{l}\in R^{l-1}\times R^{l}$ and $b\in R^{l}$ are the weight matrix and bias vector corresponding to layer $l$, and $f_{l}$ is a non linear activation function corresponding to layer $l$. Except for the input layer, all layers’ units are followed by a non linear activation function, such as the sigmoid, the hyperbolic tangent or the rectifier linear unit. When performing classification, the last layer is followed by the sigmoid activation function for binary classification or the softmax activation function for multiclass classification. A multiclass MLP with one hidden layer can is depicted in Figure 4.
@@
The MLP is commonly trained with the Backpropagation algorithm [105]. The idea is to update the model parameters in the opposite direction of the gradient, such that a loss computed from the prediction error is minimized. The algorithm is composed of a forward pass and a backward pass. In the forward pass, the input is passed through all the layers in the network, yielding a prediction. In the backward pass, the gradients of the model’s parameters are computed with respect to the loss function, making use of the chain rule. These are computed one layer at a time, iterating backward from the last layer. The computation of the gradients along with weight update are done following an optimization method, commonly using Stochastic Gradient Descent (SGD) algorithms.
@@ 4.3 Elman Recurrent Neural Network
where $h_{t}\in\mathbb{R}^{h}$ is the hidden state at time $t$, $x_{t}\in\mathbb{R}^{N}$ is the input at time $t$, $W\in\mathbb{R}^{h\times N}$, $U\in\mathbb{R}^{h\times h}$ and $b\in\mathbb{R}^{h}$ are parameters to be learned and $f$ is a non linear activation function such as the sigmoid or the hyperbolic tangent. At each timestep the RNN can produce an output:
@@
In the case of Emotion Recognition the input to the network is a sequential utterance, for example the combination of text, audio and visuals corresponding to one word being fed per timestep, and the output is the predicted emotion.
@@ 4.4 Long Short-Term Memory Networks
As in the Elman RNN, in the case of Emotion Recognition the input to the network can be a sequential utterance and the output is the predicted emotion. LSTMs are usually leveraged to model dependencies between several sequential utterances. In this case the input to the network is a sequence of utterances.
The LSTM described in the equations above does not capture information from future time steps, which is useful for example in text applications when the context of a word is made up by both previous and subsequent words in the phrase. Therefore, Bidirectional Long Short-Term Memory networks (Bi-LSTM) [33] were proposed. Bi-LSTMs consist in two LSTMs: one executing a forward pass and the other a backward pass. A sequence of hidden states is produced for each direction that are usually joined into a unique sequence of hidden states through concatenation.
@@ 4.4.1 Context-Dependent Emotion Recognition [95]
There is a strong correlation and influence in emotion distribution between sequential utterances in a conversation. Leveraging those contextual dependencies is key to the performance of a classifier. An utterance is a unit of speech which is bound by breathes or pauses [87]. In this line, the work of [95] considers interdependences among utterances. The approach consists in an architecture made of LSTMs to extract contextual features from the utterances. The model enables consecutive utterances to share information while preserving its order. The authors propose the contextual LSTM, consisting of unidirectional LSTM cells, the hidden-LSTM, in which the dense layer after the LSTM cell is omitted and the bi-directional contextual LSTM that considers information from before and after the targeted utterance. It achieved an F1-score of 54.95% on the IEMOCAP dataset.
@@ 4.5 Gated Reccurent Units
GRUs can work better than LSTMs when there is less available data.
@@ 4.5.1 DialogueRNN [78]
DialogueRNN predicts the emotion of an utterance based on the speaker, the context of preceding utterances and the emotions from preceding utterances. The model has three components all modeled by GRUs. The party state models the parties’ emotion dynamics throughout the conversation. The global state has the encoding of preceding utterances and the party state, modeling the context of the utterance. Finally, the emotion representation is based on the party state and the global state and is used to perform the final classification.
@@
It yielded an F1-score of 62.75% on the IEMOCAP dataset. The authors believe it outperforms the contextual LSTM due to better context representation.
@@ 4.6 Attention Mechanisms
The aforementioned work mentions the use of attention over surrounding emotion representaions. To motivate Attention Mechanisms (AM) [2] we now introduce Sequence to Sequence (Seq2Seq) models. Seq2Seq models are composed of a RNN encoder that takes the input sequence and outputs a context vector, the last encoder hidden state, that is fed into an RNN decoder that generates an output sequence from this vector. When the input sequence is very long the context vector does not have enough capacity to represent the information from the entire sequence. This is where AM are useful. Taking inspiration from the human visual system that attends to different parts of the space, while building its representation of the scene, the AM allow the decoder to attend to the relevant encoder hidden states. At each timestep, a context vector is obtained by a weighted sum of these hidden states, where the weights of the sum are proportional to the similarity between the current decoded hidden state and the encoded hidden states, as described in the following equations:
@@
where $h_{t}^{d}$ is the decoder hidden state and $h_{i}^{e}$ and $h_{j}^{e}$ are the encoder hidden states. The most common similarity score function is the dot product. By introducing AM in a RNN the performance of a classifier usually increases.

### v4

@@ 3.1.3 Mixed Emotions
A person can experience simultaneously several emotions. For example, one can experience joy and surprise at the same time and sometimes even opposite emotions simultaneously [3]. While the VAD model allows for an unlimited description of emotions so that a mixture of emotions can be expressed in a single VAD description, the single-label categorical approach does not account for mixed emotions. This motivates the need to classify emotions with multiple labels when using the categorical approach. This need becomes more obvious with the decreasing number of labels in a dataset, but there are also multi-label datasets with a high number of emotion labels, such as the GoEmotions dataset [13] with 28 emotions.
@@ 3.1.4 ERC Taxonomy Considerations
The existent taxonomies are usually transversally applied along all Emotion Recognition tasks. However, for ERC, it can be observed that there are coincident predominant classes in several benchmark datasets, such as Neutral and Happiness/Joy, as can be seen in the class distribution of the datasets on subsection 3.2. Within different ERC domains such as chit-chat vs task-oriented, e.g. mental health [16] and customer support [32], there are specific classes of emotions that will be more represented than others in each domain. With such a rich diversity of possible classes for this task, as it can be observed in the 28 classes of the GoEmotions dataset, the dimensional approach can encompass all classes but might make it harder to distinguish between differentiated emotional displays.
@@ 3.2 ERC Benchmark Datasets
In this subsection, different categorical, single-label datasets are presented, that were used by the works benchmarked in this survey. These works are organised in Table LABEL:ercmul. Other relevant datasets are also presented. Overall, their class imbalance is significant suggesting the use of balancing techniques that will be elaborated on in subsection 5.2. All datasets are in English, which is a limitation concerning multilingual ERC, as pointed out in section 2.
@@ 3.2.1 IEMOCAP [6]
The IEMOCAP dataset consists of 151 videos of two speakers’ dialogues, comprising 7.433 utterances with an average length of 11,6 words. Each utterrance is labelled with one of 8 emotions: happy, sad, neutral, angry, excited, frustrated fear, and disgust. Most of the works, however, benchmark their performance on the first 6 classes. The maximum class imbalance ratio is 1:3 (happy: frustrated).
Sample: “Is that, is that - is that just foam? I can’t even tell. Although, if you can’t tell, it’s probably isn’t them. It’ll probably be unmistakable, don’t you think?” Label: Excited
@@ 3.2.2 MELD [88]
The Multimodal EmotionLines Dataset (MELD) contains 13.708 utterances from 1.433 dialogues from the TV series Friends. Each utterance encompasses audio, visual, and textual modalities, with an average length of 8,0 words. Emotion categories are Ekman’s basic emotions and neutral. The maximum class imbalance ratio is 1:18 (fear: neutral).
Sample: “Are we okay now?” Label: Fear
@@ 3.2.3 EmoryNLP [135]
EmoryNLP is a text dataset extracted from transcripts of the TV show Friends. It contains 12.606 utterances, labelled with one of the six primary emotions in Willcox’s feeling wheel [119] and a neutral label. The maximum class imbalance ratio is 1:4 (sad: neutral).
Sample: “What a coincidence, I listen in my sleep.” Label: Peaceful
@@ 3.2.4 DailyDialog [62]
DailyDialog is a large text dataset built from websites used to practise English dialogue in daily life, containing 13.118 multi-turn dialogues, comprising 102.879 utterances with an average length of 14,6 words. Each utterance is labelled with one of the six Ekman’s basic emotions or other. The maximum class imbalance ratio is 1:1156 (fear:other).
Sample: “I was scared stiff of giving my first performance” Label: Disgust
Table 1 shows a comparison of the label distributions of the aforementioned datasets and Table 2 shows a comparison of the corresponding percentages.
Table 1: Label distribution comparison of different ERC datasets.
Label IEMOCAP MELD Emory NLP Daily Dialog
Neutral 1708 6436 3776 85572
Happiness/Joy 648 2308 2755 12885
Surprise - 1636 - 1823
Sadness 1084 1002 844 1150
Anger 1103 1607 1332 1022
Disgust - 361 - 353
Fear - 358 1646 74
Frustration 1849 - - -
Excitement 1041 - - -
Peace - - 1190 -
Powerful - - 1063 -
Table 2: Label percentage distribution comparison of different ERC datasets.
Label IEMOCAP MELD Emory NLP Daily Dialog
Neutral 23.0 47.0 30.0 83.2
Happiness/Joy 8.7 16.8 21.9 12.5
Surprise - 11.9 - 1.8
Sadness 14.6 7.3 6.7 1.1
Anger 14.8 11.7 10.6 1.0
Disgust - 2.6 - 0.3
Fear - 2.6 13.1 0.1
Frustration 24.9 - - -
Excitement 14.0 - - -
Peace - - 9.4 -
Powerful - - 8.4 -
@@ 3.2.5 Other Datasets
The RECOLA dataset [94] is a multimodal corpus of spontaneous human interactions collected from a collaborative task. It includes not only audio and video, but also electrocardiogram and electrodermal activity data. Emotional behaviours are annotated in terms of arousal and valence, in continuous time, and social behaviours are in terms of agreement, dominance, engagement, performance, and rapport, in discrete time.
@@ 4 Deep Learning for Emotion Recognition in Conversations
In this section, key ERC works are described, focusing on the challenges and opportunities of the conversational scenario, presented in Section 2, and on the diverse Deep Learning architectures that can be employed for the task. Descriptions of key Deep Learning models are provided along with descriptions of works that resorted to those models. The works on Emotion Recognition in Conversations using Deep Learning described in this survey were chosen to be representative of the different techniques that are employed, covering gated Neural Networks with and without bidirectionality, Memory Networks, Graph Neural Networks, Attention Mechanisms, and Transformers. It is remarkable to observe how in four years research evolved from a simple Long Short-Term Memory Network (LSTM), taking into account the context of the conversation, to complex gated and graph neural network architectures and how the invention of transformers was reflected in not only word embeddings, such as BERT, but also in new classifier architectures. Comparisons between these works are performed based on the reported weighted-F1-score without the majority class on the IEMOCAP dataset, the choice of most authors including the prominent works of Poria et al [86], Hazarika et al [31], Majumder et al [69], Ghosal et al [26], Zhong et al [144], Shen et al [98], Ghosal et al [25], Wang et al [114], Li et al [58], Shen et al [99] and Lei et al [54], since the dataset is not too imbalanced.
@@ 4.1 Convolutional Neural Networks for Feature Extraction
Convolutional Neural Networks (CNNs) are a type of feedforward neural network inspired by the biology of the visual cortex that mimics the local filtering over the input space performed by the receptive fields. In the case of text, a linear layer is applied to a sequence of words, each represented by its embeddings, separately and in the same way for each sequence. The convolutional layer can have many filters, each extracting a different feature, and is usually followed by a pooling operation. In the case of Emotion Recognition in Conversations, convolutional layers act as feature extractors, useful since strong predictors of the label can appear in different places in the input. With the advent of Pre-Trained Language Models for utterance representation, elaborated on subsection 4.10, convolutional neural networks have been less used.
4.2 The Multi-Layer Perceptron
Emotion Recognition is a classification task. The most simple neural network architecture that can be employed for such a task is the Multi-Layer Perceptron.
The Multi-Layer Perceptron (MLP) is a non-linear function approximator that can be used for classification and regression. It is a class of feed-forward artificial neural networks that is composed of at least three layers: an input layer, a hidden layer, and an output layer. Its complexity and representation power increases with the number of layers. A MLP is mathematically described by the following equation:
@@ 4.2 The Multi-Layer Perceptron
In which L is the number of layers, $W_{l}\in R^{l-1}\times R^{l}$ and $b_{l}\in\mathbb{R}^{l}$ are the weight matrix and bias vector corresponding to layer $l$, and $f_{l}$ is a non-linear activation function corresponding to layer $l$. Except for the input layer, all layers’ units are followed by a non-linear activation function, such as the sigmoid, the hyperbolic tangent, or the rectifier linear unit. When performing classification, the last layer is followed by the sigmoid activation function for binary classification or the softmax activation function for multiclass classification. A multiclass MLP with one hidden layer can is depicted in Figure 4.
@@
The MLP is commonly trained with the Backpropagation algorithm [96]. The idea is to update the model parameters in the opposite direction of the gradient, such that a loss computed from the prediction error is minimised. The algorithm is composed of a forward pass and a backward pass. In the forward pass, the input is passed through all the layers in the network, yielding a prediction. In the backward pass, the gradients of the model’s parameters are computed with respect to the loss function, making use of the chain rule. These are computed one layer at a time, iterating backward from the last layer. The computation of the gradients along with weight update are done following an optimization method, commonly using Stochastic Gradient Descent (SGD) algorithms.
@@ 4.3 Elman Recurrent Neural Network
where $h_{t}\in\mathbb{R}^{h}$ is the hidden state at time $t$, $x_{t}\in\mathbb{R}^{N}$ is the input at time $t$, $W\in\mathbb{R}^{h\times N}$, $U\in\mathbb{R}^{h\times h}$ and $b\in\mathbb{R}^{h}$ are parameters to be learned and $f$ is a non-linear activation function such as the sigmoid or the hyperbolic tangent. At each timestep the RNN can produce an output:
@@
In the case of Emotion Recognition, the input to the network is a sequential utterance, for example, the combination of text, audio, and visuals corresponding to one word being fed per timestep, and the output is the predicted emotion.
@@ 4.4 Long Short-Term Memory Networks
As in the Elman RNN, in the case of Emotion Recognition the input to the network can be a sequential utterance and the output is the predicted emotion. LSTMs are usually leveraged to model dependencies between several sequential utterances. In this case, the input to the network is a sequence of utterances.
The LSTM described in the equations above does not capture information from future time steps, which is useful for example in text applications when the context of a word is made up of both previous and subsequent words in the phrase. Therefore, Bidirectional Long Short-Term Memory Networks (Bi-LSTM) [27] were proposed. Bi-LSTMs consist of two LSTMs: one executing a forward pass and the other a backward pass. A sequence of hidden states is produced for each direction that are usually joined into a unique sequence of hidden states through concatenation.
@@ 4.4.1 Context-Dependent Emotion Recognition [86]
There is a strong correlation and influence in emotion distribution between sequential utterances in a conversation. Leveraging those contextual dependencies is key to the performance of a classifier. An utterance is a unit of speech which is bound by breathes or pauses [77]. In textual data, it is usually delimited by punctuation such as full stops, but can also be delimited by question or exclamation marks. Along these lines, the work of [86] considers interdependences among utterances. The approach consists of an architecture made of LSTMs to extract contextual features from the utterances. The model enables consecutive utterances to share information while preserving their order. The authors propose the contextual LSTM, consisting of unidirectional LSTM cells, the hidden LSTM, in which the dense layer after the LSTM cell is omitted, and the bi-directional contextual LSTM that considers information from before and after the targeted utterance. It achieved an F1-score of 54.95% on the IEMOCAP dataset.
@@ 4.5 Gated Reccurent Units
GRUs can work better than LSTMs when there is less available data since they have fewer parameters, and are less prone to overfitting.
@@ 4.5.1 DialogueRNN [69]
DialogueRNN predicts the emotion of an utterance based on the speaker, the context of preceding utterances, and the emotions from preceding utterances. The model has three components all modeled by GRUs. The party-state models the parties’ emotion dynamics throughout the conversation. The global state has the encoding of preceding utterances and the party state, and models the context of the utterance. Finally, the emotion representation is based on the party state and the global state and is used to perform the final classification.
@@
It yielded an F1-score of 62.75% on the IEMOCAP dataset. The authors argue DialogueRNN variants outperform the contextual LSTM due to better context representation.
@@ 4.6 Attention Mechanisms
The aforementioned work mentions the use of attention over surrounding emotion representations. To motivate Attention Mechanisms (AM) [1] we now introduce Sequence to Sequence (Seq2Seq) models. Seq2Seq models are composed of an RNN encoder that takes the input sequence and outputs a context vector, the last encoder hidden state, that is fed into an RNN decoder that generates an output sequence from this vector. When the input sequence is very long the context vector does not have enough capacity to represent the information from the entire sequence. This is where AMs are useful. Taking inspiration from the human visual system that attends to different parts of the space, while building its representation of the scene, the AM allows the decoder to attend to the relevant encoder hidden states. At each time step, a context vector is obtained by a weighted sum of these hidden states, where the weights of the sum are proportional to the similarity between the current decoded hidden state and the encoded hidden states, as described in the following equations:
@@
where $h_{t}^{d}$ is the decoder hidden state and $h_{i}^{e}$ and $h_{j}^{e}$ are the encoder hidden states. The most common similarity score function is the dot product. By introducing AM in an RNN the performance of a classifier usually increases.